# EEGDM: Learning EEG Representation with Latent Diffusion Model

### v1

@@ Introduction
As aforementioned, due to the inherently low SNR of original EEG signals, accurate reconstruction of them directly in the denoising process is challenging. To address this, we first enhance the SNR of the raw EEG signals using PCA. Similar to the LDMs (Rombach et al. 2022), we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals. In addition, we use the high-quality representation from the EEG encoder to guide the diffusion process. After denoising in the latent space, we apply inverse PCA to project the latent EEG representations back to the original signal space, thereby reconstructing the final EEG signals.
Diffusion Formulation
@@
To train the diffusion model with a learned reverse process covariance $\Sigma_{\theta}$, we follow the approach of Nichol et al. (2021): train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$. After the final DiT block, we use a standard linear decoder to decode the token sequence into an output noise prediction and an output diagonal covariance prediction. Both outputs have the same data shape as the input. Once $p_{\theta}$ is trained, new latent representations can be sampled by initializing $z_{t_{max}}\sim{\mathcal{N}(0,\mathbf{I})}$ and sampling $z_{t-1}\sim{p_{\theta}(z_{t-1}|z_{t})}$ via the reparameterization trick. Finally, we obtain the denoised latent representation $\widetilde{z}$, which is then projected back to the original signal space using inverse PCA to reconstruct the original EEG signals:
@@
Conditional Guidance
We incorporate the encoded EEG representation as conditional input for the diffusion model. It guides the generative process and is further leveraged for downstream tasks. To this end, we broadly consider EEG signals in this context as any collection of EEG recordings that exhibit meaningful interrelationships (e.g., semantic correlations), particularly those derived through various augmentations of source EEG signals — an approach well-established in contrastive learning frameworks (Chen et al., 2020). Specifically, each augmented EEG signal is encoded using a shared EEG encoder, and the resulting EEG representations $e^{1:m}$ are aggregated into a single vector $e$ by taking their mean, enabling the model to incorporate richer conditional information. When the conditional input is $e$, the reverse process of the diffusion model is $p_{\theta}(z_{t-1}|z_{t},e)$, where both $\epsilon_{\theta}$ and $\Sigma_{\theta}$ are conditioned on $e$. Under this setting, classifier-free guidance (Ho and Salimans 2022) can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized. According to Bayes’ rule, $p(e|z)\propto{p(z|e)/p(z)}$, and thus $\nabla_{z}\log p(e|z)\propto{\nabla_{z}\log p(z|e)-\nabla_{z}\log p(z)}$. By interpreting the output of the diffusion model as a score function, the sampling process can be guided by:
@@
Methods Model Size Balanced Accuracy ↑ AUROC ↑
@@
Table 1: Performance comparison on TUAB. The best metric is highlighted in bold, and the second-best metric is underlined.
Methods Model Size Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
@@
Table 2: Performance comparison on TUEV. The best metric is highlighted in bold, and the second-best metric is underlined.
Datasets Methods Balanced Accuracy ↑ Cohen’s Kappa ↑ Weighted F1 / AUROC ↑
@@
Table 3: Performance comparison on BCIC. The best metric is highlighted in bold, and the second-best metric is underlined.
Variants PCA EEG Augmentation BCIC-2A ↑ BCIC-2B ↑ TUAB (10%) ↑ TUEV (10%) ↑
@@
Table 4: The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
Conditional Input Setup
Experiments
Datasets
Pre-training datasets. We pre-train the proposed EEGDM on several publicly available EEG datasets covering a range of paradigms, including PhysioMI (motor imagery and execution dataset) (Goldberger et al. 2000), TSU (steady-state visual evoked potential dataset) (Wang et al. 2016), SEED (emotion recognition dataset) (Zheng and Lu 2015), and M3CV (multi-subject multi-session multi-paradigm dataset) (Huang et al. 2022).
Datasets of downstream specific tasks. (1) TUAB (abnormal detection) (Obeid and Picone 2016): A corpus of EEG recordings annotated as either normal or abnormal. (2) TUEV (event type classification) (Obeid and Picone 2016): An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG). (3) BCIC-2A (motor imagery classification) (Tangermann et al. 2012): A four-class motor imagery dataset comprising left hand, right hand, feet, and tongue movement imagination tasks. (4) BCIC-2B (motor imagery classification) (Steyrl et al. 2016): A binary motor imagery dataset containing left-hand and right-hand movement imagination tasks.
Experimental Setup
We evaluate EEGDM’s representation learning capabilities through fine-tuning on downstream tasks. Following previous work (Wang et al. 2024), we adopt standard preprocessing procedures including resampling, normalization, segmentation, and flattening. For TUAB and TUEV datasets, we use the provided train-test splits, while BCIC-2A and BCIC-2B employ Leave-One-Subject-Out (LOSO) cross-validation. The reported model sizes correspond to the downstream networks used for fine-tuning, as the generative modules are not involved in downstream tasks. We compare with state-of-the-art baselines and provide results for both EEGDM (25M parameters) and a compact variant, EEGDM-Tiny (3M parameters), which uses a smaller pre-training architecture. Detailed configurations for both model variants are provided in the supplementary material. Complete implementation details and hyperparameters for both pre-training and fine-tuning are also provided in the supplementary material.
Results and Analysis
In this section, we show that EEGDM achieves competitive performance across diverse downstream tasks, demonstrating the effectiveness of diffusion-based representation learning for EEG analysis. All experiments are conducted with three random seeds, and we report mean and standard deviation values as shown in Tables 1, 2, and 3. Additional results under different fine-tuning settings are provided in the supplementary material.
On the TUAB dataset, EEGDM achieves a balanced accuracy of 80.17%, outperforming both the previous state-of-the-art method EEGPT (79.83%) and other baselines. The AUROC score of 87.39% further demonstrates strong discriminative capability. For TUEV classification, EEGDM obtains 60.81% balanced accuracy with a Cohen’s Kappa coefficient 0.5218, achieving either (1) good balance between performance and pre-training data size, when compared to EEGPT, or (2) substantial improvements over other baselines. On motor imagery classification, EEGDM shows competitive performance, achieving 55.95% balanced accuracy on BCIC-2A and 72.19% on BCIC-2B, with the latter outperforming both EEGPT and LaBraM.
Regarding the comparison with EEGPT and LaBraM, the following detailed explanations are necessary. Following EEGPT (Wang et al. 2024), all comparisons are conducted with LaBraM-Base (Jiang, Zhao, and Lu 2024). The LaBraM results are based on the same pre-training dataset as ours, rather than their original larger dataset (which includes TUEG dataset, whose distribution is similar to the TUAB and TUEV). Remarkably, our EEGDM achieves performance comparable to or better than EEGPT despite using a smaller pre-training dataset. Even our compact EEGDM-Tiny model maintains competitive performance while using substantially fewer parameters. These results validate the effectiveness of using EEG signal generation as a self-supervised pre-training objective, demonstrating that the combination of encoder and diffusion denoiser, enhanced by PCA and channel augmentation, successfully learns high-quality EEG representations for improved downstream task performance.
Ablation Study
We conduct ablation studies to analyze the contributions of PCA and channel augmentation components, respectively, by setting up four variant versions as shown in Table 4, in which the average balanced accuracy on three random seeds is reported. For BCIC-2A and BCIC-2B, we report the results using the LOSO evaluation protocol. Variant D consistently achieves the best performance across all settings, demonstrating that both PCA and EEG augmentation are important for our model. Furthermore, for TUAB and TUEV, we randomly selected 10% of the training set for fine-tuning and tested it on the entire test set (see the last two columns in Table 4); we also use this setting in Figure 3. The results show that the advantages of incorporating PCA and EEG augmentation remain evident even when small-scale datasets are used for fine-tuning.
We also investigate the impact of the number of PCA components on downstream task performance. Figure 3 presents results on TUEV and BCIC-2B with varying numbers of PCA components, which reveal a nonlinear relationship between the number of PCA components and downstream task performance. In particular, insufficient number of components may result in large information loss, whereas an excessive number introduces noise contamination, both of which degrade model performance. To balance these trade-offs, we empirically select an intermediate value of 20 components, which yields optimal performance in our experiments.
To evaluate generation quality, we use the diffusion model in EEGDM under the same pre-training configuration, and compute Pearson correlation coefficients between the generated and original signals in both time and frequency domains. When PCA is applied, the correlation coefficients improve significantly to 0.36 (time domain) and 0.88 (frequency domain), compared to lower values of 0.003 and 0.83 without PCA. These results demonstrate that integrating PCA substantially improves the quality of EEG data generation, which further facilitates better representation learning.
@@
Figure 3: Performance comparison with different numbers of PCA components on TUEV and BCIC-2B datasets.
Visualization
To evaluate the quality of learned representations, we visualize the EEG data embeddings using t-SNE (Van der Maaten and Hinton 2008) as presented in Figure 4. We randomly sample EEG signals from the test sets of four downstream tasks in a class-balanced manner and extract their representations using our pre-trained encoder. The t-SNE visualizations exhibit well-separated clusters across all datasets, indicating that our EEGDM learns discriminative representations capable of capturing task-relevant patterns. This clear clustering structure further demonstrates the EEGDM’s ability to learn generic yet informative EEG representations that can be successfully adapted to diverse downstream tasks. Beyond representation learning, EEGDM also achieves high-quality EEG signal reconstruction, as evidenced by visualizations of reconstructed signals and their power spectral density (PSD) analyses provided in the supplementary material.
(a) TUAB
(b) TUEV
(c) BCIC-2A
(d) BCIC-2B
@@
In summary, our experimental results highlight three key advantages of EEGDM: (1) High-quality EEG reconstruction, as validated by the visualizations in supplementary material; (2) Effective representation learning, evidenced by distinct class separation in t-SNE visualizations and consistent performance improvements across tasks; (3) Competitive performance across diverse downstream tasks using modest pre-training data sizes, particularly given that EEGDM matches or exceeds EEGPT’s performance while using smaller pre-training datasets.
Conclusion

### v2

@@ III-C EEG Data Generation by Diffusion Model
As aforementioned, due to the inherently low SNR of original EEG signals, accurate reconstruction of them directly in the denoising process is challenging. To address this, we first enhance the SNR of the raw EEG signals using PCA. Similar to the LDMs [rombach2022high], we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals. In addition, we use the high-quality representation from the EEG encoder to guide the diffusion process. After denoising in the latent space, we apply inverse PCA to project the latent EEG representations back to the original signal space, thereby reconstructing the final EEG signals.
III-C1 Diffusion Formulation
@@
To train the diffusion model with a learned reverse process covariance $\Sigma_{\theta}$, we follow the approach of Nichol et al. [nichol2021improved]: train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$. After the final DiT block, we use a standard linear decoder to decode the token sequence into an output noise prediction and an output diagonal covariance prediction. Both outputs have the same data shape as the input. Once $p_{\theta}$ is trained, new latent representations can be sampled by initializing $z_{t_{max}}\sim{\mathcal{N}(0,\mathbf{I})}$ and sampling $z_{t-1}\sim{p_{\theta}(z_{t-1}|z_{t})}$ via the reparameterization trick. Finally, we obtain the denoised latent representation $\widetilde{z}$, which is then projected back to the original signal space using inverse PCA to reconstruct the original EEG signals:
@@
III-C2 Conditional Guidance
We incorporate the encoded EEG representation as conditional input for the diffusion model. It guides the generative process and is further leveraged for downstream tasks. To this end, we broadly consider EEG signals in this context as any collection of EEG recordings that exhibit meaningful interrelationships (e.g., semantic correlations), particularly those derived through various augmentations of source EEG signals — an approach well-established in contrastive learning frameworks (Chen et al., 2020). Specifically, each augmented EEG signal is encoded using a shared EEG encoder, and the resulting EEG representations $e^{1:m}$ are aggregated into a single vector $e$ by taking their mean, enabling the model to incorporate richer conditional information. When the conditional input is $e$, the reverse process of the diffusion model is $p_{\theta}(z_{t-1}|z_{t},e)$, where both $\epsilon_{\theta}$ and $\Sigma_{\theta}$ are conditioned on $e$. Under this setting, classifier-free guidance [ho2022classifier] can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized. According to Bayes’ rule, $p(e|z)\propto{p(z|e)/p(z)}$, and thus $\nabla_{z}\log p(e|z)\propto{\nabla_{z}\log p(z|e)-\nabla_{z}\log p(z)}$. By interpreting the output of the diffusion model as a score function, the sampling process can be guided by:
@@ III-C2 Conditional Guidance
III-C3 Conditional Input Setup
IV Experiment
IV-A Datasets
Pre-training datasets. We utilize the Temple University Hospital EEG Corpus (TUEG) [obeid2016temple] to pre-train the proposed EEGDM. TUEG is a very large publicly available clinical EEG dataset containing clinical-grade recordings collected in real-world hospital environments. It has over 40 different channel configurations and varying temporal durations of recordings. The majority of the recordings are acquired at a sampling rate of 256 Hz. TUEG exhibits substantial heterogeneity in patient demographics, channel configurations, and recording equipment, making it particularly suitable for pre-training our model.
Datasets of downstream specific tasks. (1) TUEV (event type classification) [harati2015improved]: An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG). The EEG signals are recorded at 23 channels and 256 Hz sampling rate. (2) TUAB (abnormal detection) [lopez2015automated]: A corpus of EEG recordings annotated as either normal or abnormal. Similar to TUEV, the EEG signals of TUAB are recorded at 23 channels and 256 Hz sampling rate. (3) SHU-MI (motor imagery classification) [ma2022large]: An EEG motor imagery dataset designed for decoding imagined movements. It features 32 EEG channels sampled at 250 Hz, including two types of motor imagery tasks: left hand and right hand. (4) CL-Drive (cognitive load assessment) [angkan2024multimodal]: It is a multimodal dataset designed for cognitive load assessment, which includes EEG signals. The EEG data are recorded using four channels at a sampling frequency of 256 Hz. This study explores the ternary classification of cognitive load into high, medium, and low levels. (5) Seed-VII (emotion recognition) [jiang2024seed]: It is a large-scale multimodal emotion dataset. Seed-VII includes 62-channel EEG and eye-movement data from 20 healthy subjects across seven emotion categories (disgust, fear, sad, neutral, happy, anger, surprise). The preprocessed EEG signals were downsampled to 200 Hz. (6) DMER (emotion distribution prediction) [yang2024multimodal]: It is a multimodal emotion recognition corpus that includes EEG signals along with other modalities (GSR, PPG, and frontal facial videos). In each trial, participants watched video clips and subsequently completed a self-assessment using a ten-item version of the Positive Affect and Negative Affect Schedule (PANAS) [mackinnon1999short], reporting the intensity of ten emotional adjectives. Each emotion rating ranges from 1 (not at all) to 5 (extremely), and the resulting scores are transformed into a 10-dimensional emotion distribution. The EEG recordings comprise 18 channels sampled at 300 Hz.
It is worth noting that we primarily select downstream datasets from the past five years, a strategic shift from the benchmarks used in works like CBraMod [wang2024cbramod] or EEGPT [wangeegpt]. This decision stems from the need to evaluate the model against the fidelity and complexity of modern neuroscience standards. Older datasets often lack the high spatial density and signal quality required to stress-test a foundation model, whereas recent paradigms — ranging from naturalistic stimuli to multimodal retrieval — better assess the capacity for generalizable, high-level cognitive decoding. However, to ensure fair comparison with established baselines, we make exceptions for the TUAB and TUEV datasets. Despite falling outside the five-year window, these two widely adopted benchmarks are retained as standardized anchors, allowing us to validate our model’s performance on classical clinical tasks alongside its capabilities on modern, complex scenarios.
IV-B Experimental Setup
For the pre-training dataset TUEG, following prior work [wang2024cbramod], we select 19 common EEG channels. All EEG signals are resampled to 200 Hz and segmented into non-overlapping 30-second samples (19 channels × 6,000 time points). Each EEG patch is set to a temporal length of 1 second, resulting in 570 patches per sample. For all downstream datasets, we maintain consistency with our pre-training data by resampling EEG signals to 200 Hz and setting each EEG patch duration to 1 second.
For the downstream datasets TUEV and TUAB, consistent with previous studies [wangeegpt, jiang2024large, wang2024cbramod], we utilize 16 common bipolar montage channels from the international 10-20 system, with sample durations of 5 seconds and 10 seconds, respectively. We adopt the train-test splits provided by these datasets. For the CL-Drive dataset, we use the channel configuration and sample duration (10 seconds) provided by the dataset and apply 10-fold subject-wise cross-validation. For the SHU-MI, DMER, and SEED-VII datasets, we use the channel configurations provided by each dataset and segment the signals into 4-second samples. We employ a subject-independent split strategy for these datasets. We fine-tune EEGPT [wangeegpt], LaBraM-Base [jiang2024large] (the only variant with available pre-trained weights), and CBraMod [wang2024cbramod] based on their publicly available code and pre-trained weights, except when their experimental results are already reported in the original papers.
For the EEG generator (diffusion module) and encoder, we adopt the DiT and ViT architectures, respectively. Following common practices in generative modeling literature, we maintain an exponential moving average (EMA) of model weights with a decay rate of 0.9999 during pre-training. The specific hyperparameters for pre-training are detailed in Table I.
We evaluate the representation learning capabilities of EEGDM through fine-tuning on downstream tasks. Specifically, we predict corresponding labels using a linear predictor based on the outputs of the EEG encoder. Following the common evaluation protocols, we present results under the fine-tuning setting, where the predictor is jointly trained with the EEG encoder (i.e., the pre-trained encoder is not frozen). We employ cross-entropy loss for downstream classification tasks and KL divergence loss for downstream distribution prediction tasks. The hyperparameters for fine-tuning EEGDM on downstream tasks are provided in Table II. We conduct a series of quantitative and qualitative experiments to evaluate the proposed model, demonstrating its strong representation capabilities across diverse downstream datasets.
TABLE I: Hyperparameters for EEGDM pre-training.
Hyperparameter Value
Optimization Batch size 32
Peak learning rate 5e-4
Minimal learning rate 5e-5
Learning rate scheduler OneCycleLR
Weight decay 0.05
EMA decay rate 0.9999
Gradient clipping norm 1
Optimizer AdamW
${\beta}_{1}$ 0.9
${\beta}_{2}$ 0.999
EEG Generator Layers N 8
Hidden size 512
Heads 8
Normalization type LayerNorm
EEG Encoder Layers 8
Hidden size 512
MLP size 2048
Heads 8
Sampling Classifier-free guidance 9
Diffusion training steps 1000
TABLE II: Hyperparameters for downstream fine-tuning. ($\star$) Depends on the dataset.
Hyperparameter Value
Batch size 64
Peak learning rate 5e-4
Minimal learning rate 5e-5
Learning rate scheduler OneCycleLR
Optimizer AdamW
${\beta}_{1}$ 0.9
${\beta}_{2}$ 0.999
Weight decay 0.05
Total epochs 30, 150
Warmup Epochs 6, 30
IV-C Results and Analysis
In this section, we show that EEGDM achieves competitive performance across diverse downstream tasks, demonstrating the effectiveness of diffusion-based representation learning for EEG analysis. All experiments are conducted with multiple random seeds, and we report the mean and standard deviation of evaluation metrics. The reported model size corresponds to the downstream networks used for fine-tuning, as the generative module is not involved in downstream tasks. Results are compared against baselines, as shown in Tables III, IV, V and VI. Compared to other approaches, EEGDM consistently maintains either the best or second-best performance in key metrics (balanced accuracy or KL divergence).
TABLE III: Performance comparison on TUEV. The best metric is highlighted in bold, and the second-best metric is underlined.
Methods Params Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
@@ IV-C Results and Analysis
[…]
TABLE IV: Performance comparison on TUAB. The best metric is highlighted in bold, and the second-best metric is underlined.
Methods Params Balanced Accuracy ↑ AUROC ↑ AUC-PR ↑
@@
[…]
TABLE V: Performance comparison on DMER. The best metric is highlighted in bold, and the second-best metric is underlined.
Methods Chebyshev ↓ Clark ↓ Canberra ↓ KL ↓ Cosine ↑ Intersection ↑
[…]
TABLE VI: The results of different methods on various datasets. The best metric is highlighted in bold, and the second-best metric is underlined.
Datasets Methods Balanced Accuracy ↑ Weighted F1 / AUROC ↑ Cohen’s Kappa / AUC-PR ↑
@@
TABLE VII: The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
Variants PCA EEG Augmentation TUEV TUAB (10%)
[…]
On the TUEV dataset, EEGDM achieves a balanced accuracy of 68.19%, outperforming both the previous state-of-the-art method CBraMod (66.71%) and other baselines. This superiority highlights the exceptional discriminative power of the representations learned by the diffusion model for classifying different event types in EEG signals.
Table IV shows that EEGDM achieves a balanced accuracy of 81.80% and an AUC-PR of 90.85% on the TUAB dataset. Securing the second-best performance on those metrics, EEGDM overall demonstrates comparable performance with a competitive margin. The consistent performance across both balanced accuracy and AUC-PR metrics suggests that EEGDM learns robust representations that generalize well for binary abnormal detection tasks, maintaining high discriminative power even in challenging clinical scenarios.
Table V presents the results for the DMER dataset, which uses distributional metrics. EEGDM achieves best performance across various divergence and similarity metrics, with the exception of the Clark metric, where it ranks second. The overall high performance on DMER, a task requiring the prediction of a continuous emotion distribution rather than a discrete class, strongly suggests that the diffusion-based pre-training learns rich and fine-grained features that preserve the complex, subtle, and continuous emotional states embedded in the EEG signals.
Table VI further validates the generalizability and robustness of EEGDM across diverse BCI tasks. Specifically, on the SHU-MI dataset, EEGDM ranks second in terms of balanced accuracy (63.24%) and AUC-PR (70.70%), while achieving the best performance in AUROC (69.96%), demonstrating competitive performance compared with CBraMod. For the CL-Drive dataset, EEGDM shows the best balanced accuracy and the best Cohen’s Kappa, demonstrating superior cognitive load feature extraction capabilities. On the Seed-VII dataset, EEGDM continues its strong performance with the best performance in balanced accuracy and ranks second in weighted F1 and Cohen’s Kappa, further validating its effectiveness in emotion-related tasks.
These comprehensive results across diverse datasets validate the effectiveness of using EEG signal generation as a self-supervised pre-training objective, demonstrating that the combination of encoder and diffusion denoiser, enhanced by PCA and channel augmentation, successfully learns high-quality EEG representations for improved downstream task performance.
(a) TUEV
(b) TUAB
@@
Figure 3: Performance comparison with different numbers of PCA components on TUEV and TUAB datasets.
IV-D Ablation Study
We conduct ablation studies to analyze the contributions of PCA and channel augmentation components, respectively, by setting up four variant versions as shown in Table VII, in which the average balanced accuracy on three random seeds is reported. Variant D consistently achieves the best performance across all settings, demonstrating that PCA and EEG augmentation are important for our model. However, we find that using PCA or EEG augmentation alone is either detrimental to or has minimal impact on our model’s performance. This may be attributed to the fact that PCA reduces the dimensionality of the source signal, making the model more sensitive to the diversity of conditional information, while data augmentation provides sufficient diversity to support learning in this reduced-dimensional space. Consequently, the PCA and EEG augmentation operations achieve complementary effects in terms of generalization performance and accuracy. Furthermore, for TUAB, we randomly selected 10% of the training set for fine-tuning and tested it on the entire test set (see the last column in Table VII); we also use this setting in Fig. 3. The results show that the advantages of incorporating PCA and EEG augmentation remain evident even when small-scale datasets are used for fine-tuning.
We also investigate the impact of the number of PCA components on downstream task performance. Fig. 3 presents results on TUEV and TUAB with varying numbers of PCA components, which reveal a nonlinear relationship between the number of PCA components and downstream task performance. In particular, insufficient number of components may result in large information loss, whereas an excessive number introduces noise contamination, both of which degrade model performance. To balance these trade-offs, we empirically select an intermediate value of 20 components, which yields optimal performance in our experiments.
(a) TUEV
(b) TUAB
@@ IV-D Ablation Study
Figure 5: Visualization of reconstructed EEG signals and their PSD.
IV-E Visualization
To evaluate the quality of learned representations, we visualize the EEG data embeddings using t-SNE [van2008visualizing] as presented in Fig. 4. We randomly sample EEG signals from the test sets of two downstream tasks in a class-balanced manner and extract their representations using our pre-trained encoder. The t-SNE visualizations exhibit well-separated clusters across all datasets, indicating that our EEGDM learns discriminative representations capable of capturing task-relevant patterns. This clear clustering structure further demonstrates the EEGDM’s ability to learn generic yet informative EEG representations that can be successfully adapted to diverse downstream tasks.
Beyond representation learning, EEGDM also achieves high-quality EEG signal reconstruction, as evidenced by visualizations of reconstructed signals and their power spectral density (PSD) analyses. As shown in Fig. 5, the reconstructed EEG signals on the pre-training dataset, as well as their frequency-domain representations, closely match the ground truth. While minor discrepancies remain at the fine-grained level, the global temporal structure and frequency characteristics are well preserved. These results indicate that the proposed model can generate realistic EEG signals and effectively capture high-level semantic information from EEG data.
In summary, our experimental results highlight three key advantages of EEGDM: (1) High-quality EEG reconstruction, as validated by signal reconstruction visualization; (2) Effective representation learning, as demonstrated by distinct class separation in t-SNE visualization; and (3) Competitive performance across diverse downstream tasks, as evidenced by consistent performance improvements across tasks.
V Conclusion